# Robust Blockwise Random Pivoting:

### v3

@@ 6.2 Accuracy and Efficiency of ID
In Figures 2, 3, 4, 5, and 6, we compare the empirical accuracy (left) and efficiency (right) of different ID algorithms on data matrices described in Section 6.1, with a focus on those demonstrating robust skeleton complexities on the adversarial input Example 4.1 (i.e. algorithms in Figure 1 (right)).
Figure 2: Relative interpolation error and runtime of ID algorithms with robust skeleton complexities on the GMM adversarial input (Example 4.1). Recall from Figure 1 that sampling methods, especially squared-norm sampling, suffer from significantly higher skeleton complexities than sequential/robust blockwise random pivoting and sketchy pivoting on GMM. For demonstration, we omit squared-norm sampling from the plot.
Taking an integrated view of ID including skeleton selection and interpolation matrix construction, we consider the following relative ID error777 All error measurements are normalized with $\left\|\mathbf{X}\right\|_{F}^{2}$ to relative errors. and runtime measurements:
@@
    For error-reveling algorithms (i.e. except for sketchy pivoting and sampling)888 In Figures 2, 3, 4, 5, and 6, we also omit the residual errors of the sequential error-reveling algorithms SRP and CPQR (which overlap with the corresponding interpolation errors) for conciseness of demonstration. , we plot the residual errors (denoted as residual) from their on-the-fly skeletonization error estimates and the runtimes for skeleton selection (e.g. $\left\|\mathbf{d}^{(t)}\right\|_{1}$ in Algorithm 4) only.
@@
    For exactly ID-revealing algorithms (i.e. except for sketchy pivoting and sampling), we plot the runtimes and interpolation errors $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)$ with $\mathbf{W}$ from Algorithm 5 (ID), which are equal to the corresponding skeletonization errors $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)=\mathcal{E}_{\mathbf{X}}\left(S\right)$.
@@
    For approximately ID-revealing algorithms (i.e. sketchy pivoting), we plot the runtimes and interpolation errors $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)$ with $\mathbf{W}$ from both Algorithm 5 (ID) and Algorithm 6 (OSID) with small multiplicative oversampling $l=2k$, where both $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)>\mathcal{E}_{\mathbf{X}}\left(S\right)$, and Algorithm 6 tends to be more expensive.
  4. (d)
    For non-ID-revealing algorithms (i.e. sampling), without explicitly specifying in the legends, we plot the runtimes and interpolation errors $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)=\mathcal{E}_{\mathbf{X}}\left(S\right)$ with the optimal $\mathbf{W}$ from solving the exact least square problem in Equation 8.
The legends in Figures 2, 3, 4, 5, and 6 are formatted in terms of “skeleton selection algorithm”-“ID error”. The abbreviations of different skeleton selection algorithms are summarized in Table 2, whereas the ID error measurements are described as above.
@@
Remark 6.1 (Choice of block size).
The choice of block size $b$ controls the trade-off between parallelizability and error-revealing property. To be precise, an unsuitably small $b$ lessens the efficiency gained from moving flops to matrix-matrix multiplications, whereas the larger $b$ leads to higher overestimation of the target rank. In the experiments, we use $b=30$ or $40$ to exploit the parallelizability of matrix-matrix multiplications without considerably compromising the error-revealing property.
@@
For the adversarial GMM dataset in Figure 2, it is worth mentioning that the sharp transition of RBGP at $\left|S\right|\approx 100$ is because GMM contains $100$ clusters, each with distinct squared-norms. Recall that blockwise greedy pivoting picks points with the maximum norms without randomness. When $\left|S\right|\leq 100$, each block drains points from the same clusters with the maximum residual norms first, most of which are redundant points and are later excluded by the robust blockwise filtering. This leads to the inefficiency of RBGP when $\left|S\right|\leq 100$. Such inefficiency of RBGP in comparison to RBRP again illustrates the superiority of random pivoting over greedy pivoting, i.e. the merit of randomness.
Table 3: GPU runtime of the relatively efficient/accurate ID algorithms in Figure 2 on the synthetic adversarial GMM dataset (Example 4.1) of size $10^{5}\times 10^{3}$, with $k=100$ cluster centers. Experiments were performed on an Nvidia V100 GPU. Our code relied on the magma_dgeqp3_gpu routine from the MAGMA library  [17] for CPQR and the cuBLAS library [38] for other matrix operations.
@@
From experiments on the (non-adversarial) natural and synthetic datasets in Figures 3, 4, 5, and 6, we observe the following (by algorithms):
  • The robust blockwise random pivoting (RBRP) enjoys (one of) the best skeleton complexities, with the skeletonization error estimates aligning perfectly with the interpolation errors (left), demonstrating the error-revealing and exactly ID-revealing abilities of RBRP. The skeleton selection process of RBRP tends to take slightly longer than that of sketchy pivoting (SkLUPP/SkCPQR) with ID (which is not error-revealing and suffers from much larger interpolation errors than RBRP). However, RBRP is generally faster than sketchy pivoting (SkLUPP/SkCPQR) with small multiplicative oversampling (OSID), while enjoying lower interpolation error. Overall, RBRP is shown to provide the best combination of accuracy and efficiency when the explicit construction of ID (in contrast to the skeleton subset only) is inquired.
  • The sketchy pivoting algorithms (SkLUPP/SkCPQR) also enjoy (one of) the best skeleton complexities, despite the lack of error-revealing ability and being approximately ID-revealing (i.e. suffering from large interpolation errors). That is, for skeleton selection only (without asking for the interpolation matrix), sketchy pivoting provides the most efficient selection of close-to-the-best skeleton subsets in practice (cf. Figure 1). Moreover, our experiments verify the conclusion in [16] that SkLUPP tends to be more efficient than SkCPQR in selecting skeletons without compromising the skeletonization error.
  • The inherently sequential algorithms, including SRP and CPQR, are highly competitive in terms of accuracy while being considerably slower than the parallelizable ones.
@@ 7 Discussion
In this work, we focus on fast and accurate ID algorithms from five perspectives that measure the empirical performance systematically: (a) skeleton complexity, (b) asymptotic complexity, (c) parallelizability, (d) error-revealing property, and (e) ID-revealing property. With a careful exploration of various existing ID algorithms through the lens of adaptiveness and/or randomness, we reemphasize the effectiveness of combining adaptiveness and randomness in solving ID problems, while unveiling a critical missing piece in the family of existing algorithms that achieves all five of the aforementioned desired properties. To close the gap, we propose robust blockwise random pivoting (RBRP)—a parallelizable, error-revealing, and exactly ID-revealing algorithm that enjoys among-the-best skeleton and asymptotic complexities in practice compared to the existing algorithms and demonstrates promising robustness to adversarial inputs. Through numerical experiments, we illustrate the empirical accuracy, efficiency, and robustness of RBRP on a broad spectrum of natural and synthetic datasets.
@@ Acknowledgments
The authors wish to thank Yifan Chen, Ethan Epperly, Kevin Miller, Chris-topher Musco, Rachel Ward, and Robert Webber for enlightening discussions.

### v5

@@ 6.2 Accuracy and Efficiency of ID
In Figures 2, 3, 4, 5, 6, and 7, we compare accuracy (left) and efficiency (right) of different ID algorithms on data matrices in Section 6.1.
Figure 2: Relative interpolation error and total runtime of ID algorithms on the GMM adversarial input (Example 4.1). Recall from Figure 1 that in lack of adaptiveness, methods like SqNorm and BRP tend to suffer from suboptimal skeleton complexities on GMM.
Taking an integrated view of ID including skeleton selection and interpolation matrix construction, we consider the following relative ID error666 All error measurements are normalized with $\left\|\mathbf{X}\right\|_{F}^{2}$ to relative errors. and runtime measurements:
@@
    For ID-revealing algorithms (i.e. except for sketchy pivoting and sampling), we plot the runtimes and interpolation errors $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)$ with $\mathbf{W}$ from Algorithm 3 (ID), which are equal to the corresponding skeletonization errors $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)=\mathcal{E}_{\mathbf{X}}\left(S\right)$.
@@
    For sketchy pivoting, we plot the runtimes and interpolation errors $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)$ with $\mathbf{W}$ from both Algorithm 3 (ID) and Algorithm 4 (OSID) with small multiplicative oversampling $l=3k$. Notice that both ID and OSID have $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)>\mathcal{E}_{\mathbf{X}}\left(S\right)$, while OSID in Algorithm 4 tends to be more expensive because of oversampling.
@@
    For squared-norm sampling, without explicitly specifying in the legends, we plot the runtimes and interpolation errors $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)=\mathcal{E}_{\mathbf{X}}\left(S\right)$ with the optimal $\mathbf{W}$ from solving the exact least-square problem in (6).
The legends in Figures 2, 3, 4, 5, 6, and 7 are formatted in terms of “skeleton selection algorithm”-“ID error”. The abbreviations of different skeleton selection algorithms are summarized in Table 2, whereas the ID error measurements are described above.
@@
Figure 7: Relative interpolation error and runtime of ID algorithms on Helmholtz.
Remark 6.3 (Choice of block size).
The choice of block size $b$ controls the trade-off between hardware efficiency and rank-adaptiveness. To be precise, an unsuitably small $b$ lessens the efficiency gained from moving flops to matrix-matrix multiplications, whereas the larger $b$ leads to a higher overestimation of the target rank. In the experiments, we use $b=30$ or $40$ to exploit the hardware efficiency of matrix-matrix multiplications without considerably compromising the rank-adaptiveness.
For the adversarial GMM dataset in Figure 2, it is worth mentioning that the sharp transition of RBGP at $\left|S\right|\approx 100$ is because GMM contains $100$ clusters, each with distinct squared-norms. Recall that blockwise greedy pivoting picks points with the maximum norms without randomness. When $\left|S\right|\leqslant 100$, each block drains points from the same clusters with the maximum residual norms first, most of which are redundant points and are later excluded by the robust blockwise filtering. This leads to the inefficiency of RBGP when $\left|S\right|\leqslant 100$. Such inefficiency of RBGP in comparison to RBRP again illustrates the superiority of random pivoting over greedy pivoting, i.e. the merit of randomness.
Table 3: GPU runtime of the relatively efficient/accurate ID algorithms in Figure 2 on an adversarial GMM matrix (Example 4.1) of size $10^{5}\times 10^{3}$, with $k=100$ cluster centers. Experiments were performed on an Nvidia V100 GPU. Our code relied on the magma_dgeqp3_gpu routine from the MAGMA library  [26] for CPQR and the cuBLAS library [55] for other matrix operations.
@@
From the (non-adversarial) datasets in Figures 3, 4, 5, 6, and 7, we observe the following:
  • RBRP enjoys (one of) the best skeleton complexities. In terms of efficiency, RBRP tends to be slower than sketchy pivoting (SkLUPP) with ID (which is neither rank-adaptive nor ID-revealing). However, RBRP is usually faster than sketchy pivoting with OSID, while enjoying lower interpolation error. Overall, RBRP is shown to provide the best combination of accuracy and efficiency when the explicit construction of ID (in contrast to the skeleton subset only) is inquired.
  • Sketchy pivoting (SkLUPP) also enjoy (one of) the best skeleton complexities, despite the lack of rank-adaptiveness and ID-revealing property (i.e. suffering from large interpolation errors). That is, for skeleton selection only (without asking for the interpolation matrix), sketchy pivoting provides the most efficient selection of close-to-the-best skeleton subsets in practice (cf. Figure 1).
  • The sequential algorithms SRP and CPQR are highly competitive in terms of accuracy while being considerably slower than the hardware-efficient blockwise algorithms.
@@
  • When using SqNorm as an ID algorithm, constructing the interpolation matrix by solving (6) explicitly can slow down the overall ID algorithm significantly to the extent that the runtimes are similar to those of RBRP. Moreover, sampling methods like SqNorm and DPP lack rank-adaptiveness and ID-revealing ability.
@@ 7 Discussion
In this work, we focus on fast and accurate ID algorithms from five perspectives that measure the empirical performance systematically: (a) skeleton complexity, (b) asymptotic complexity, (c) hardware efficiency, (d) rank-adaptiveness, and (e) ID-revealing property. With a careful exploration of various existing ID algorithms through the lens of adaptiveness and/or randomness, we reemphasize the effectiveness of combining adaptiveness and randomness in solving ID problems, while unveiling a critical missing piece in the family of existing algorithms that achieves all five of the aforementioned desired properties. To close the gap, we propose robust blockwise random pivoting (RBRP)—a hardware-efficient, rank-adaptive, and ID-revealing algorithm that enjoys among-the-best skeleton and asymptotic complexities in practice compared to the existing algorithms and demonstrates promising robustness to adversarial inputs. Through numerical experiments, we illustrate the empirical accuracy, efficiency, and robustness of RBRP on a broad spectrum of natural and synthetic datasets.
@@ Acknowledgments
The authors wish to thank Yifan Chen, Ethan Epperly, Kevin Miller, Christopher Musco, Rachel Ward, and Robert Webber, as well as the anonymous referees and editors, for enlightening discussions and constructive feedback.